# HaHeAE: Learning Generalisable Joint Representations of Human Hand and Head Movements in Extended Reality

### v1

@@ V-A Identifying Interpretable Hand-head Clusters
(b) Activity: Learn course while sitting. The head turns to the left; both arms are bent and have no large movements.
@@
(d) Activity: Take a tape. The head is facing slightly upward; the left hand has a greater range of motion than the right hand.
Figure 3: Representative hand and head movements (unit: metres) for each of the four largest clusters and their semantics on the EgoBody dataset. The red and blue lines indicate the trajectories of left and right hands, respectively, while the black arrows denote head orientations. The colours of the lines and arrows are gradually deepened over time.
We observed that the largest cluster (a) mainly consists of noticeable horizontal movements of both hands across different activities such as instruct to act, teach course, perform, and dance. Furthermore, we found that cluster (b) and cluster (c) contain very limited hand-head movements at different locations, suggesting that the user mainly stood or sat still and barely moved. The samples forming these clusters are mainly generated during the activities that require few hand and head movements such as learn course while sitting and casually chat while standing. In these scenarios, users mainly stay still and may only use slight hand or head movements (e.g., gestures or nodding) as the body language during communications [47], in line with the patterns illustrated by these two clusters. In addition, cluster (d) demonstrates moving one hand and the other hand mainly staying still. These samples were generated from activities requiring only one hand to interact with objects such as take a tape or pick up a cup. The above results demonstrate that our method can be applied to identify common patterns of hand-head movements and thus has great potential to facilitate the analysis of hand-head signals.
@@ V-B Generating Variable Hand-head Movements
Collecting large-scale human behavioural data in extended reality serves as the basis for understanding human behaviours [7, 8, 21] and developing data-driven models [4, 23, 5]. However, human behavioural data collection is usually costly and time-consuming and raises serious concerns about privacy and security [48, 4, 49]. To address these challenges, one common solution is to augment existing datasets instead of collecting new data [12, 13]. Our method can be used for data augmentation by generating hand-head movements with controllable variations. Specifically, we first added random noise to the stochastic representation $E_{sto}$ learned by our method using the formula:
$E_{var}=E_{sto}+\beta z,$ (4)
where $z$ is a random noise sampled from $N(0,I)$ and has the same size as $E_{sto}$, $\beta$ is a fixed constant used as a weighting factor to control the variation, and $E_{var}$ is the altered stochastic representation. We then used our decoder (see Figure 1) to generate hand-head movements from the semantic and altered stochastic representations. This way, our method can generate hand-head movement variations that differ in details while preserving the core semantic information. Figure 4 shows some generation examples on the EgoBody dataset. We can see that the generated hand-head movements are similar to the original one in overall moving trend, with notable differences in moving trajectory. In addition, we also find that the difference between the generated movement and the original one becomes larger with the increasing weighting factor, demonstrating the weighting factor’s effectiveness in controlling variations. The above results illustrate the usefulness of our method as an effective means to generate variable hand-head movements, which has significant potential for enriching and augmenting existing hand-head movement datasets.
Figure 4: Hand and head movements (unit: metres) generated by adding weighted Gaussian noises to the stochastic representations learned by our method on the EgoBody dataset. Our method can generate variable hand-head movements that differ in details while preserving core semantics. The weighting factor can effectively control the degree of variations.
@@ VI-A On Performance
In this work, we proposed HaHeAE – the first self-supervised method to learn generalisable joint representations of human hand and head movements in extended reality. We evaluated our method in the challenging out-of-domain setting, i.e., we trained it on one XR dataset and tested it on two unseen datasets across different users, activities, and XR environments. We showed that our method significantly outperforms other methods in hand and head reconstruction performance, achieving improvements of up to 74.0% (see Table I). We also demonstrated that our method has strong generalisation capability for different users, activities, and XR environments (see Table I). We further evaluated the learned semantic representations on two sample downstream tasks relevant to XR applications: understanding users and their interaction context. As shown in Table II, the semantic representations learned by our method consistently outperform those learned by other methods across different tasks and different datasets, achieving an improvement of up to 5.2% (55.7% vs. 50.5% on EgoBody) in terms of activity recognition accuracy. Taken together, these results are highly promising and, for the first time, demonstrate the feasibility of learning generalisable joint representations of hand and head movements in a self-supervised fashion. That is, without requiring additional human annotations that are usually costly, cumbersome, and time-consuming to collect [15]. As such, our work opens up an exciting new research direction for hand-head behaviour modelling in extended reality.
@@ VI-B On the Method
Even more interestingly, by disentangling the hand-head joint representations into a semantic and a stochastic part, our method enables novel use cases with significant potential for XR applications. Specifically, the semantic representation can be used to analyse hand-head behaviours in XR by identifying interpretable hand-head clusters within the original signals (see Figure 3) while the stochastic representation can be used to generate hand-head movements with controllable variations (see Figure 4). Given these results, our method contributes to the explainability of human hand and head behaviours in XR and can serve as an analysis or generation tool to facilitate future research in human hand-head behaviour modelling and help develop future human-aware intelligent XR systems. Furthermore, our method also provides meaningful insights into learning disentangled representations of human behaviours in extended reality.
@@ VI-D Limitations and Future Work
Despite all these advances, we also identified several limitations that we plan to address in future work. First, to learn generalisable representations across different XR environments, we used relative coordinates with the origin set to the position of the head. However, in doing so we neglected the translational movement of the head and only encoded the head’s rotational movement in our representation. In the future, we plan to explore how to also include translational components of the head movements into the learned latent embeddings. In addition, our method only encodes the hand and head movements without considering the context information, such as the XR scene content, audio signals, or the specific activities that the users are performing. Learning joint representations from more modalities beyond hand and head movements is an interesting avenue of future work and would provide a more comprehensive understanding of human behaviours in extended reality. Furthermore, our method takes observed hand and head signals as input but such input may be incomplete in real applications due to tracking errors. How to deal with incomplete hand-head data also requires further study. Finally, in the future, we would also like to explore more applications of our hand-head joint representations to assist XR researchers in analysing and understanding human behaviours.

### v3

@@ V-A Identifying Interpretable Hand-head Clusters
(b) Activity: Learn course while sitting. The head is facing slightly upward; both arms are bent and have no large movements.
@@
(d) Activity: Take a tape. The head is facing forward; the left hand has a greater range of motion than the right hand.
Figure 3: Representative hand and head movements for each of the four largest clusters and their semantics on the EgoBody dataset. The red and blue lines indicate the trajectories of left and right hands, respectively, while the black arrows denote head orientations. The colours of the lines and arrows are gradually deepened over time.
We observed that the largest cluster (a) mainly consists of noticeable horizontal movements of both hands across different activities such as instruct to act or perform. Furthermore, we found that cluster (b) and cluster (c) contain very limited hand-head movements at different locations, suggesting that the user mainly stood or sat still and barely moved. The samples forming these clusters are mainly generated during the activities that require few hand and head movements such as learn course while sitting or casually chat while standing. In these scenarios, users mainly stay still and may only use slight hand or head movements (e.g., gestures or nodding) as the body language during communications [50], in line with the patterns illustrated by these two clusters. In addition, cluster (d) demonstrates moving one hand and the other hand mainly staying still. These samples were generated from activities requiring only one hand to interact with objects such as take a tape or pick up a cup. The above results demonstrate that our method can be applied to identify common patterns of hand-head movements and thus has great potential to facilitate the analysis of hand-head signals. Furthermore, we also compared the clustering performance of our method with that of the baseline methods and validated the superiority of our method (see supplementary material for more details).
@@ V-B Generating Variable Hand-head Movements
Collecting large-scale human behavioural data in extended reality serves as the basis for understanding human behaviours [7, 8, 21] and developing data-driven models [4, 23, 5]. However, human behavioural data collection is usually costly and time-consuming and raises serious concerns about privacy and security [51, 4, 52]. To address these challenges, one common solution is to augment existing datasets instead of collecting new data [12, 13]. Our method can be used for data augmentation by generating hand-head movements with controllable randomness from the original signals. Specifically, we first added random noise to the stochastic representation $E_{sto}$ learned by our method using the formula:
$E_{var}=E_{sto}+\beta\cdot z,$ (4)
where $z$ is a random noise sampled from $N(0,I)$ and has the same size as $E_{sto}$, $\beta$ is a fixed constant used as a weighting factor to control the randomness, and $E_{var}$ is the altered stochastic representation. We then used our decoder (see Figure 1) to generate hand-head movements from the semantic and altered stochastic representations. This way, our method can generate hand-head movement variations that differ in details while preserving the core semantic information. Figure 4 shows some generation examples on the EgoBody dataset. We can see that the generated hand-head movements are similar to the original one in overall moving trend, with notable differences in moving trajectory. To further explore our method’s applicability, we replaced our stochastic representation with random noise $z$ to generate hand-head signals. The results in Figure 4 (column $z$) showed that our method can generate variable and realistic hand-head movements from random noise. The above results illustrate the usefulness of our method as an effective means to generate variable and realistic hand-head movements, which has significant potential for enriching and augmenting existing hand-head movement datasets.
(a) Original
(b) $E_{sto}$
(c) $E_{sto}$ + $0.1\cdot z$
(d) $E_{sto}$ + $0.3\cdot z$
(e) $z$
Figure 4: Hand and head movements generated from altered stochastic representations and random noise on the EgoBody dataset. Our method can generate variable and realistic hand-head movements from altered stochastic representations and random noise.
@@ VI-A On Performance
In this work, we proposed HaHeAE – the first self-supervised method to learn generalisable joint representations of human hand and head movements in extended reality. We evaluated our method in the challenging out-of-domain setting, i.e., we trained it on one XR dataset and tested it on two unseen datasets across different users, activities, and XR environments. We showed that our method significantly outperforms other methods in hand and head reconstruction performance, achieving improvements of up to 74.1% (see Table I). We also demonstrated that our method has strong generalisation capability for different users, activities, and XR environments (see Table I). We further evaluated the learned semantic representations on two sample downstream tasks relevant to XR applications: understanding users and their interaction context. As shown in Table II, the semantic representations learned by our method consistently outperform those learned by other methods across different tasks and different datasets, achieving an improvement of up to 5.2% (55.7% vs. 50.5% on EgoBody) in terms of activity recognition accuracy. Taken together, these results are highly promising and, for the first time, demonstrate the feasibility of learning generalisable joint representations of hand and head movements in a self-supervised fashion. That is, without requiring additional human annotations that are usually costly, cumbersome, and time-consuming to collect [15]. As such, our work opens up an exciting new research direction for hand-head behaviour modelling in extended reality.
@@ VI-B On the Method
Even more interestingly, by disentangling the hand-head joint representations into a semantic and a stochastic part, our method enables novel use cases with significant potential for XR applications. Specifically, the semantic representation can be used to analyse hand-head behaviours in XR by identifying interpretable hand-head clusters within the original signals (see Figure 3) while the stochastic representation can be used to generate hand-head movements with controllable randomness (see Figure 4). Given these results, our method contributes to the explainability of human hand and head behaviours in XR and can serve as an analysis or generation tool to facilitate future research in human hand-head behaviour modelling and help develop future human-aware intelligent XR systems. Furthermore, our method also provides meaningful insights into learning disentangled representations of human behaviours in extended reality.
@@ VI-D Limitations and Future Work
Despite all these advances, we also identified several limitations that we plan to address in future work. First, to learn generalisable representations across different XR environments, we used relative coordinates with the origin set to the position of the head. However, in doing so we neglected the translational movement of the head and only encoded the head’s rotational movement in our representation. In the future, we plan to explore how to also include translational components of the head movements into the learned latent embeddings. In addition, we trained our method only on the EgoBody dataset and directly tested it on the ADT and GIMO datasets. In future work we would like to train and test our method on more datasets to further evaluate its generalisability. Finally, in the future, we would also like to explore more applications of our hand-head joint representations to assist XR researchers in analysing and understanding human behaviours.